14. Two identical firms compete as a Cournot duopoly. The demand they face is P = 100 - 2Q. The cost function for each firm is C(Q) = 4Q. In equilibrium, the deadweight loss is: (a) $128, (b)$256, (c) $384, (d) $512, (e) none of them are true..

15. Two identical firms compete as a Cournot duopoly. The demand they face is P = 100 - 2Q. The cost function for each firm is C(Q) = 4Q. The equilibrium output of each firm is: (a) 8, (b) 16, (c) 32, (d) 36, (e) none of them are true.

16. Two identical firms compete as a Cournot duopoly. The demand they face is P = 100 - 2Q. The cost function for each firm is C(Q) = 4Q. Each firm earns equilibrium profits of: (a) $1,024, (b) $2,048, (c)$4,096, (d) $512 (e) none of them are true

17. Two identical firms compete as a Cournot duopoly. The demand they face is P = 100 - 2Q. The cost function for each firm is C(Q) = 4Q. In equilibrium, the deadweight loss is: (a)$128, (b)$256, (c)$384, (d)$512, (e) none of them are true.

18. Two firms compete as a Stackelberg duopoly. The demand they face is P = 100 - 3Q. The cost function for each firm is C(Q) = 4Q. The outputs of the two firms are:

(a).QL = 16; QF = 8.

(b).QL = 24; QF = 12.

(c).QL = 12; QF = 8.

(d).QL = 20; QF = 15.

(e). None of them are true.

19. Two firms compete as a Stackelberg duopoly. The demand they face is P = 100 - 3Q. The cost function for each firm is C(Q) = 4Q. The profits of the two firms are:

20. Suppose a manager is interested in implementing third-degree price discrimination. The manager knows that the price elasticity of demand for Group 1 is -2 and the price elasticity of demand for Group 2 is -1.2. Based on this information alone we can conclude that the price charged to Group 2 will be

(a) the same as the price charged to Group 1.

(b) lower than the price charged to Group 1.

(c) higher than the price charged to Group 1.

(d) there is insufficient information to determine whether Group 2 will have a higher,

lower or the same price as Group 1.

(e) none of them are true.
Answer #1

Answering only first four MCQ are mandatory by the HomeworkLib

Q14) option B) 256

15) option B) 16

​​​​​

16) option D)

Each firm profit = (P-MC)q

As Q = 32, So P = 100-64 = 36

π = (36-4)* 16

= 512

17) option B ) 256

18) OPTION A)

in stackelberg game, if P = a - bQ

& MC = c

Then output of leader = (a-c)/2b

= (100-4)/2*3

= 96/6 = 16

Output of follower firm, = (a-c)/4b

= 96/4*3

= 8